solve If the sequence a1, a2, a3, … an form an A. P.,  then a1^2  - a2^2  + a3^2  - … + a2n – 1^2  - a2n^2  =
 

Grade:11

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

a^2 – b^2 = (a+b)(a-b)

here a-b = d ie common difference which can be taken common

so the summation of the sequence is

( d)(Sn)

Regards